### 01 Feb 2024 v1

@@ Appendix A Msss matyices
$\Delta_{67}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-(\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu)^{2}\tan\beta}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (84)
@@
$\Delta_{77}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (85)
@@
$\Delta_{6(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-(\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b})^{2}\mu\tan\beta\lambda_{i}v_{u}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (86)
@@
$\Delta_{7(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b})^{2}\mu\lambda_{i}v_{u}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (87)
@@
$\Delta_{(7+i)(7+j)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u})^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}).$ (88)
@@
$M_{SP}^{2}=\left(\begin{array}[]{ccccc}M_{h_{d}\sigma_{d}}^{2}&M_{h_{d}\sigma_{u}}^{2}&M_{h_{d}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{h_{d}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{h_{d}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{h_{u}\sigma_{d}}^{2}&M_{h_{u}\sigma_{u}}^{2}&M_{h_{u}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{h_{u}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{h_{u}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{(\tilde{\nu}_{1}^{c})^{\Re}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{(\tilde{\nu}_{2}^{c})^{\Re}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{(\tilde{\nu}_{3}^{c})^{\Re}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ \end{array}\right)$ (94)
@@
$M_{h_{d}(\tilde{\nu}_{i}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=-\Im(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}+\Delta_{1(7+i)},$ (99)
$M_{h_{u}(\tilde{\nu}_{i}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=-\Im(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}+\Delta_{2(7+i)},$ (100)
$M_{(\tilde{\nu}_{i}^{c})^{\Re}\sigma_{d}}^{2}\!\!\!\!\!$ $=-\Im(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}+\Delta_{(2+i)6},$ (101)
$M_{(\tilde{\nu}_{i}^{c})^{\Re}\sigma_{u}}^{2}\!\!\!\!\!$ $=\Im(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}+\Delta_{(2+i)7},$ (102)
$M_{(\tilde{\nu}_{i}^{c})^{\Re}(\tilde{\nu}_{i}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=2\Im(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}v_{u}$ (103)
$-(A_{\kappa}\kappa)_{i}\Re(e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})\upsilon_{\nu_{i}^{c}}+\Delta_{(2+i)(7+i)},$
$M_{(\tilde{\nu}_{i}^{c})^{\Re}(\tilde{\nu}_{j}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=\Delta_{(2+i)(7+i)}.$ (104)
@@
$\Delta_{16}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{|\mu|^{2}\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (105)
×$\times{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
@@
$\Delta_{17}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{|\mu|^{2}\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (106)
×$\times{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\tan\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
@@
$\Delta_{26}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (107)
@@
×$\times\Bigg{\{}\ln{m_{\tilde{t}_{1}}^{2}\over m_{\tilde{t}_{2}}^{2}}+{{|A_{t}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
@@
$\Delta_{27}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (108)
@@
×$\times\Bigg{\{}\ln{m_{\tilde{t}_{1}}^{2}\over m_{\tilde{t}_{2}}^{2}}+{{|A_{t}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
@@
$\Delta_{1(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{|\mu|^{2}\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (109)
×$\times{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
@@
$\Delta_{2(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (110)
@@
×$\times\Bigg{\{}\ln{m_{\tilde{t}_{1}}^{2}\over m_{\tilde{t}_{2}}^{2}}+{{|A_{t}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$

### 06 Mar 2024 v2

@@ Acknowledgements.
$\Delta_{67}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-(\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu)^{2}\tan\beta}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (84)
@@
$\Delta_{77}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (85)
@@
$\Delta_{6(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-(\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b})^{2}\mu\tan\beta\lambda_{i}v_{u}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (86)
@@
$\Delta_{7(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b})^{2}\mu\lambda_{i}v_{u}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (87)
@@
$\Delta_{(7+i)(7+j)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u})^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}).$ (88)
@@
$M_{SP}^{2}=\left(\begin{array}[]{ccccc}M_{h_{d}\sigma_{d}}^{2}&M_{h_{d}\sigma_{u}}^{2}&M_{h_{d}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{h_{d}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{h_{d}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ M_{h_{u}\sigma_{d}}^{2}&M_{h_{u}\sigma_{u}}^{2}&M_{h_{u}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{h_{u}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{h_{u}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ \end{array}\right)$ (94)
@@
$M_{h_{d}(\tilde{\nu}_{i}^{c})^{\text{Im}}}^{2}\!\!\!\!\!$ $=-\text{Im}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}+\Delta_{1(7+i)},$ (99)
$M_{h_{u}(\tilde{\nu}_{i}^{c})^{\text{Im}}}^{2}\!\!\!\!\!$ $=-\text{Im}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}+\Delta_{2(7+i)},$ (100)
$M_{(\tilde{\nu}_{i}^{c})^{\text{Re}}\sigma_{d}}^{2}\!\!\!\!\!$ $=-\text{Im}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}+\Delta_{(2+i)6},$ (101)
$M_{(\tilde{\nu}_{i}^{c})^{\text{Re}}\sigma_{u}}^{2}\!\!\!\!\!$ $=\text{Im}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}+\Delta_{(2+i)7},$ (102)
$M_{(\tilde{\nu}_{i}^{c})^{\text{Re}}(\tilde{\nu}_{i}^{c})^{\text{Im}}}^{2}\!\!\!\!\!$ $=2\text{Im}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}v_{u}$ (103)
$-(A_{\kappa}\kappa)_{i}\text{Re}(e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})\upsilon_{\nu_{i}^{c}}+\Delta_{(2+i)(7+i)},$
$M_{(\tilde{\nu}_{i}^{c})^{\text{Re}}(\tilde{\nu}_{j}^{c})^{\text{Im}}}^{2}\!\!\!\!\!$ $=\Delta_{(2+i)(7+i)}.$ (104)
@@
$\Delta_{16}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{|\mu|^{2}\cot\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (105)
×$\times{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
@@
$\Delta_{17}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{|\mu|^{2}\cot\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (106)
×$\times{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\tan\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
@@
$\Delta_{26}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (107)
@@
×$\times\Bigg{\{}\ln{m_{\tilde{t}_{1}}^{2}\over m_{\tilde{t}_{2}}^{2}}+{{|A_{t}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
@@
$\Delta_{27}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (108)
@@
×$\times\Bigg{\{}\ln{m_{\tilde{t}_{1}}^{2}\over m_{\tilde{t}_{2}}^{2}}+{{|A_{t}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
@@
$\Delta_{1(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{|\mu|^{2}\cot\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (109)
×$\times{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
@@
$\Delta_{2(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (110)
@@
×$\times\Bigg{\{}\ln{m_{\tilde{t}_{1}}^{2}\over m_{\tilde{t}_{2}}^{2}}+{{|A_{t}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$